# AI sustains higher strategic tension than humans in chess

### v1

@@ Abstract
Strategic decision-making involves managing the tension between immediate opportunities and long-term objectives. We study this trade-off in chess by characterizing and comparing dynamics between human vs. human and AI vs. AI games. We propose a network-based metric of piece-to-piece interaction to quantify the ongoing strategic tension on the board. Its evolution in games reveals that the most competitive AI players sustain higher levels of strategic tension for longer durations than elite human players. Cumulative tension varies with algorithmic complexity for AI and correspondingly in human-played games increases abruptly with expertise at about 1600 Elo and again at 2300 Elo. The profiles reveal different approaches. Highly competitive AI tolerates interconnected positions balanced between offensive and defensive tactics over long periods. Human play, in contrast, limits tension and game complexity, which may reflect cognitive limitations and adaptive strategies. The difference may have implications for AI usage in complex, strategic environments.
Chess — Artificial Intelligence — Complex Networks — Strategic Tension — Cognitive Science
@@
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term wins must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 37, 3], of which war [43, 4] and conflicts [39, 11, 5] are examples. In the more general case, organisms and individuals learn or, over evolutionary time scales, adapt to changing ecological niches such as during development (e.g., neural pruning [44]), foraging (e.g., depending on climate [28]), and social change (e.g., reorganization of social hierarchy [21]). These can be characterized as strategic choices involving an implicit trade-off between overfitting the local environment and maintaining the flexibility to adjust to unforeseen environmental conditions [19]. A complete measurement of the strategic environment, costs, and even the outcome is not readily available in many cases. The game of chess, in contrast, presents a long-studied and bounded example of competitive strategic interaction, where the decisions made by the players and their consequences for the game outcome are completely measurable.
@@
As a result, the game of chess has been the model system for investigating strategy. It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [33, 38, 8, 29]. Chess is a competitive game between two white and black players who take turns making moves with a single piece on a board with 8x8 squares, as shown in Figure 1. Both players start with 16 pieces, and the game aims to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves that prohibits exhaustive calculation, and human players must thus make heuristic decisions about the potential long-term value of moves [38]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [16, 17, 15]. In contrast, modern AI, particularly deep reinforcement learning systems like AlphaZero [35] and Leela Chess Zero [20], display superhuman performance by developing strategies through self-play. Their strategies differ significantly from human approaches [36, 34, 22]. Alongside these, widely used engines like Stockfish [40] rely primarily on brute-force search with handcrafted evaluation functions and deep search trees, incorporating machine learning only marginally. In principle, this suggests that the nature of the trade-offs that they need to make are distinct from human players, but there is little characterization of properties that distinguish their gameplay.
@@
Here, we compare humans to AI play by analyzing the network structure of piece-to-piece interactions [12] over many chess games. We consider thousands of games across a broad range of scenarios, including games with players across all skill levels and from official AI tournaments between competitive chess engines such as Leela Chess Zero and Stockfish, held between 2019 and 2024. In these tournaments, the engines operate adaptively, allocating varying amounts of time to different moves. To these, we add games generated locally with limited computational power, in which Stockfish is run at fixed search depths and thus with constant computational effort per move.
@@
We find that the complexity of the game, as measured by “strategic tension,” follows a regular curve, whose features correspond to the different stages of the game. The curve is characteristically different for human plays compared with AI plays. Our observations indicate that humans and AI behave differently in complex strategic environments and raise questions about the repercussions of AI managing competitive settings.
@@ I Defining Tension in Chess
Given the state of a chessboard as in Figure 1a, we consider the set of threatening and defensive relationships at each ply (half-move). The threatening piece would be one able to capture an opposing piece in one move. The defending piece would threaten any piece that would capture the piece it is defending. This is the corresponding weighted network in Figure 1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
@@
  • Defense link: We define a defense link from $p_{i}$ to $p_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally capture $p_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ in a single move. Kings can defend other pieces, but cannot be defended under this definition. Defense links reveal the layers of protection that influence whether an attack can be safely carried out or needs to be postponed.
@@
In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board. Attack and defense links can have a weight of 1 or 2, depending on whether the corresponding conditions are satisfied by one or both pieces involved. In contrast, control links always have a weight of 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
@@
Figure 1: One example of tension network. (a) Chess board at the ply 44 during a game between the two grandmasters Magnus Carlsen (white) and Alexey Sarana (black) in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark and light circles, depending on whether they are black or white. Attack links are orange, when originating from white pieces, and red, when originating from black. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Vacant squares under control are shown as small brownish nodes. The chosen ply represents the highest tension reached in each game according to the maximum eigenvalue of the adjacency matrix.
@@
Among the many ways of characterizing the interaction network, we are particularly interested in how the complexity of the game evolves, captured in features like the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces that form a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which is attacking the rook on b4). of interacting pieces, as described above. Importantly, we interpret this complexity not just in terms of structural richness, but also in terms of strategic fragility, the potential for cascading consequences following a suboptimal move. In dense, highly connected configurations, even a small error may propagate across the network, leading to substantial shifts in advantage. To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, for which we borrow an analogous, technical term from chess analysis, “strategic tension.” This quantity summarizes how tightly coupled the interaction network is at any given moment, and has been shown to reflect critical structural dynamics [27]. For simplicity, we consider the undirected tension network, which ensures a unique and real-valued largest eigenvalue.
@@
The maximum eigenvalue reflects network functionality, including its dynamical properties and specific structural features [6]. This connection can be understood through the lens of topological entropy, a concept from dynamical systems that captures the complexity and unpredictability of paths in a network [1]. In this context, the spectral radius—the largest eigenvalue—reflects the network’s inherent “diversity” of possible trajectories. A higher value indicates greater dynamical complexity, mirroring the idea of tension as a measure of uncertainty or potential instability in the system. Importantly, the spectral radius is bounded between the average and the maximum node degree, i.e., $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, which emphasizes its role as a structural proxy capturing both global connectivity and local heterogeneity [41]. Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) as a measure of game complexity at the current state of the board (more details in the Supplementary Materials).
Given a network with $N=64$ nodes and an adjacency matrix whose entries are at most 2, the highest possible maximum eigenvalue is T=2(N1)=126T=2(N-1)=126, which occurs when the network is fully connected and each edge has weight 2. However, in the context of chess games, the adjacency matrices are quite sparse, and the resulting tension is significantly lower than this theoretical upper bound. Due to the constraints imposed by legal chess moves, such a fully connected weighted network is not achievable in practice. The example shown in Figure 1, which corresponds to the ply with the highest tension, has a tension value of approximately $T\simeq 6.65$. For reference, the maximum tension that can be reached by two players iteratively choosing legal moves that most increase tension is around $T\simeq 10.3$ at ply 52, with limited room for further increases (see Supplementary Materials).
Result 1: AIs tolerate more tension
@@
We show the typical evolution of the strategic tension throughout many games in Figure 2. We can distinguish four main phases. The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces start approaching each other, and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 50, represents the peak of complexity, where the board remains highly interconnected and the tension stays elevated. The third exhibits a simplification of the tension network, as exchanges and reductions gradually lower the overall tension to about ply 120. The final and fourth is characterized by a low but relatively stable tension, as the remaining pieces maneuver without significantly altering the balance. The third and fourth phases are strongly influenced both by the overall game length and by the eventual outcome. Thus, the tension captures the overall contours of the game, as discussed in the theory of chess play [10], and indicates that the maximum eigenvalue is a meaningful proxy for game complexity and the potential for cascading effects. We do not report tension values beyond ply 150, as the available statistics become increasingly sparse—especially for games played between humans, which typically conclude earlier (Figure 2b).
We compare the tension profile with the Stockfish evaluation of the same games, which quantifies the game board’s state. We find that sharp changes in evaluation often occur at or shortly after peaks in tension [2], supporting the idea that these moments mark critical points where a single suboptimal move can trigger significant consequences (see Supplementary Materials).
@@
Following the peak, the rate at which the tension decreases depends on several factors. Among these, whether a game ends in a draw or not plays only a marginal role compared to the distinction between human-played and engine-played games. Generally, draws are significantly more common among chess engines, and tension lasts longer (Figure 2c). Moreover, the peak tension typically occurs at a later ply and tends to be lower on average compared to human games. Within human games, we observe that those ending in a draw tend to exhibit lower tension. This is likely due to many drawn games’ more cautious and symmetrical nature, where early simplifications and fewer sharp imbalances keep the tension low. In contrast, tension values are not as clearly differentiated between drawn and decisive outcomes in games played between AIs.
Result 2: Complementary Metrics Beyond Tension
@@
Tension captures general and coarse features of the game. For example, AI maintains a higher number of pieces on average, reflecting less simplification early on. The interaction networks derived from engine games display more links and loops, indicating a richer and more resilient web of mutual threats, defenses, and controls. That the maximum average number of loops is, on average, higher in human games, as with the higher peak in tension, indicates similarity between the two metrics, which are discussed further in “ Supplementary Materials.”
@@
Figure 3: Comparison of structural properties across human and AI games. We analyze the set of games shown in Figure 2. Two measures are considered: the average standard deviation of the degrees associated with chess pieces, normalized by the mean of the degrees, and the average tension per piece. Red is for AI and blue for human games.
As with tension, engine play in the endgame shows more variation in how pieces are connected to one another. We measure this with the noise coefficient of the degree distribution over chess pieces (standard deviation over the mean degree), which follows a similar trend in both human and AI games up until the final stages, where they start to diverge at approximately ply 80 (Figure 3a). There, AI exhibits greater variability in the degree distribution. This indicates that, on average, human players tend to distribute interactions more evenly across their pieces, while AI tends to stress interactions on specific pieces. This pattern aligns with the higher tension values in the endgame because a more heterogeneous distribution of degrees facilitates the emergence of highly connected hubs, increasing the adjacency matrix’s maximum eigenvalue.
@@
This difference in tension between AI and human play, however, is inverted when normalized by the number of chess pieces remaining on the board, $\lambda_{1}/n_{t}$. As shown in Figure 3b, it remains consistently higher across all plies in human games. While human play shows an approximately linear increase in tension per piece throughout the game, AI games display a sublinear progression that appears to plateau around ply 80. This suggests that while AI maintains “tactical pressure” more persistently during late-game phases compared to human play, it does so with significantly more pieces remaining on the board. To quantify better this effect, we also investigated the correlation between tension values and the number of chess pieces present on the board (see Supplementary Materials).
While we have thus far treated edges in the graph equally, they are different. In particular, the surplus of attack links over defense links, or Attack-Defense Balance (Figure 4), highlights how the early phase of the game tends to be more defense-oriented. As the game progresses, however, attack links gradually outnumber defense links. This shift is noticeably more pronounced in human games compared to those played by AI.
@@
Figure 4: Balance between attack and defense links across human and AI games. We analyze the set of games shown in Figure 2. For both human and AI games, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively.
This behavior depends on different strategic styles, especially for drawn AI games, which a lower Attack-Defense Balance characterizes. Human players often seek direct and decisive operations, favoring sharper, more aggressive transitions when tactical chances arise. In contrast, AI maintains a more balanced configuration in the long run, resulting in wider piece coordination, latent threats, and structurally embedded defenses. In draws, AI games prefer solidity over risk, with extended defensive networks and fewer explicit attacks. This pattern also aligns with the tension trends shown in Figure 2, where tension remains high and decreases slowly in drawn AI games. Since defense links can only form in response to attacks, a high tension level in the later phases of the game likely reflects a dense but balanced interaction, where attacks tend to be answered by appropriate defenses. Extending the parallel with the tension profile, we observe that in this case, it is human games that reach a higher peak of tension, whereas in the case of the Attack-Defense Balance, AI games exhibit a more pronounced minimum. This comparison shows that tension aligns with a balance between attack and defense.
@@
Overall, we find that strategic tension aligns with basic metrics about gameplay, such as the number of pieces and loops in the interaction graph. Still, it is likewise helpful to compare it with the balance across pieces on the board and how the pieces are arranged in formation. The picture is that AIs maintain more pieces, a more balanced network of attack and defense connections, and display more variability in the tension across the pieces they retain.
@@
Result 3: Tension increases with skill
@@
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for human games (blue dots) and AI games between Stockfish and Leela from official TCEC tournaments (red dot), all associated with estimated Elo ratings. Each dot represents the average across 1,200 games. The blue dot associated with the highest Elo rating and the red dot were derived from the data analyzed in Figure 2. (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot.

### v2

@@ Abstract
Strategic decision-making requires balancing immediate opportunities against long-term objectives: a tension fundamental to competitive environments. We investigate this trade-off in chess by analyzing the dynamics of human and AI gameplay through a network-based metric that quantifies piece-to-piece interactions. Our analysis reveals that elite AI players sustain substantially higher levels of strategic tension for longer durations than top human grandmasters. We find that cumulative tension scales with algorithmic complexity in AI systems and increases linearly with skill level (Elo rating) in human play. Longer time controls are associated with higher tension in human games, reflecting the additional strategic complexity players can manage with more thinking time. The temporal profiles reveal contrasting approaches: highly competitive AI systems tolerate densely interconnected positions that balance offensive and defensive tactics over extended periods, while human players systematically limit tension and game complexity. These differences have broader implications for understanding how artificial and biological systems navigate complex strategic environments and for the deployment of AI in high-stakes competitive scenarios.
Chess — Artificial Intelligence — Complex Networks — Strategic Tension
@@
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term gains must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 40, 3], of which war [weinsteinRebellionPolitics2007, bohorquezCommonEcology2009] and conflicts [smithLogicAnimal1973, dedeoInductiveGame2010a, brushConflictsInterest2018] are poignant examples. In the more general case, organisms and individuals learn such as during development (e.g., neural pruning [woudeNoGains2019]), foraging (e.g., depending on climate [piteraDailyForaging2018]), and social change (e.g., reorganization of social hierarchy [28]) or, over evolutionary time scales, “adapt” to changing ecological niches [24]. In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33]. In many real-world settings, a complete measurement of the strategic environment, costs, and outcomes is not available. Chess, by contrast, is a game of perfect information, providing a well-studied and bounded setting.
@@
As a result, chess has served as a model system for investigating strategy. It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32]. Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig. 1. Both players start with 16 pieces, and the objective of the game is to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves, which prohibits exhaustive calculation. As a result, human players must make heuristic decisions about the possible long-term value of moves [39]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [21, 25, 20]. In contrast, modern AI exhibits superhuman performance through two main strategies. One class involves deep reinforcement learning systems, such as AlphaZero and Leela Chess Zero, which develop strategies through self-play using deep convolutional neural networks. Alongside these, engines like Stockfish combine massive brute-force search trees with NNUE (Efficiently Updatable Neural Networks). While Stockfish originally relied on handcrafted evaluation functions, modern versions utilize machine learning to train their evaluation layers, blending classical search efficiency with neural network-based intuition. These differences suggest that the nature of the trade-offs that AI engines make is distinct from those made by human players. Still, there is little characterization of the properties that distinguish gameplay.
@@
Here, we compare humans to AI play by analyzing the network structure of piece-to-piece interactions [16] over many chess games. We consider thousands of games across a broad range of scenarios, including games with players of all skill levels and official AI tournaments between competitive chess engines, such as Leela Chess Zero and Stockfish, held between 2019 and 2024. In these tournaments, the engines operate adaptively, allocating varying amounts of time to different moves. In addition to these tournament games, we include simulations in which Stockfish is run at fixed search depths. This serves as a controlled baseline, allowing us to evaluate piece-piece interaction networks under uniform computational constraints.
@@
We find that the complexity of the game, as measured by “strategic tension,” follows, on average, a regular curve, whose features correspond to the different stages of the game. The curve is characteristically different for human players compared with AI players. Our observations suggest that humans and AI exhibit distinct behaviors in complex strategic environments, raising questions about the consequences of delegating competitive real-life scenarios to AI.
@@ I Defining Tension in Chess
In chess, tension refers to a state of unresolved tactical interaction where multiple pieces from both sides occupy squares from which they can capture or be captured by one another. Unlike a direct exchange of pieces, which resolves the interaction and simplifies the board state, tension represents a “latent” complexity. It persists as long as players maintain these mutual threats without acting upon them, forcing both competitors to constantly evaluate the shifting consequences of a potential explosion of captures. By formalizing these threats into a network of offensive and defensive relationships, we can quantify this tension as the density and structure of piece-to-piece interactions.
Given the state of a chessboard as in Fig. 1a, we consider the set of offensive and defensive relationships at each ply (half-move). An attacking piece can capture an opposing piece in a single move. The defending piece would threaten any piece that would capture the piece it is defending. This is the corresponding weighted network in Fig. 1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
@@
  • Defense link: We define a defense link from $p_{i}$ to $p_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally capture $p_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ in a single move. Kings can defend other pieces, but cannot be defended under this definition. This distinction arises because the King is never captured in chess; hence, the legal mechanism of a reciprocal capture (or “re-capture”) cannot apply to it. Defense links reveal the layers of protection that influence whether an attack can be safely carried out or needs to be postponed.
@@
In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board. Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved. In contrast, control links always have weight 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our case studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
@@
Figure 1: One example of tension network. (a) Chessboard at the ply 44 during a game between the two grandmasters, Magnus Carlsen (white) and Alexey Sarana (black), in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark and light circles, depending on whether they are black or white. Attack links are displayed in orange when originating from white pieces and in red when originating from black. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Small, brownish nodes represent vacant squares under control. The chosen ply represents the highest tension reached during the game according to the maximum eigenvalue of the adjacency matrix.
@@
Among the many ways to characterize the interaction network, we focus on how game complexity evolves, reflected in features such as the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces forming a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which attacks the rook on b4).. We interpret this complexity both in terms of structural richness and strategic fragility, namely the potential for cascading consequences following a suboptimal move. In dense, highly connected configurations, even a small error may propagate across the network, shifting advantage substantially. To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, adopting the chess-inspired term strategic tension. This quantity summarizes how tightly coupled the interaction network is at any moment and reflects critical structural dynamics [31]. For simplicity, we consider the undirected tension network, treating tension as a reciprocal property between pieces.
@@
The maximum eigenvalue reflects network functionality, encompassing both its structural features and dynamical properties [8]. This can be understood via topological entropy, which captures the complexity and unpredictability of paths in a network [1]: the spectral radius summarizes the network’s “diversity” of possible trajectories. Higher $\lambda_{1}$ indicates greater dynamical complexity, consistent with tension as a measure of uncertainty or potential instability. Importantly, $\lambda_{1}$ is bounded between the average and maximum node degree, $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, linking global connectivity with local heterogeneity [41]. Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) at the current state of the board (see the Supplementary Materials for details).
RESULT 1: humans peak, AI sustains STRATEGIC TENSION
@@
We illustrate the typical evolution of the strategic tension across multiple games in Fig. 2. We can distinguish four main phases. The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces approach each other, and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated. The third phase exhibits a gradual simplification of the tension network, as piece exchange and loss steadily lower the overall tension. Finally, the games reach a relatively stable value of tension, but at this point available statistics become increasingly sparse, especially for games played between humans that typically end earlier (Fig. 2b and 2c). Thus, the dynamics of tension capture the overall contours of the game, including the opening buildup, the combinatorially complex midgame, and the relaxation leading up to the endgame as discussed in the theory of chess play [13].
Despite the overall similarities in the typical trajectory of strategic tension between human and AI play, there are notable differences. Maximum tension is higher in human-played games than in engine-played games. Up to approximately ply 40, human players typically follow established opening repertoires from theoretical databases, which often lead to well-studied positions with numerous direct piece interactions, including attacks, defenses, and immediate threats [15, 6]. These theory-driven positions generate the highest points of tension as players transition to original play in the midgame. Once players leave the realm of memorized theory and must rely on their own calculation and evaluation, they tend to simplify positions. Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations. In contrast, state-of-the-art engines tend to maintain more distributed, positional advantages that unfold gradually over multiple moves, resulting in a slower decay of tension.
A secondary difference emerges in how strategic tension declines in drawn vs. decisive games. Drawn games between humans show a rapid and substantial fall in tension relative to decisive games. This agrees with the faster decrease in the number of pieces for drawn games between human players (see Fig. 10 in the Supplementary Materials). AI games show a contrasting pattern. Drawn games show significantly more (not fewer) pieces on the board relative to decisive games in the endgame and tension does not differentiate strongly.
II Result 2: Strategic tension reflects, does not predict game result
To relate the structural complexity of the interaction network to game outcomes, we compare the evolution of strategic tension with the algorithmic evaluation provided by Stockfish (see Fig. 8 in the Supplementary Materials). Inspecting the evaluation score, we identify bifurcation points in game trajectories, typically around ply 30 (move 15), where decisive and drawn games begin to diverge for human players during the opening-to-midgame transition. A similar behavior is observed in AI games, where a local maximum or inflection point in the evaluation profile precedes the divergence of outcomes. Notably, in all games, this divergence occurs before strategic tension reaches its absolute peak. This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws. In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established while the board is still increasing in complexity.
The ensuing reduction in tension reflects a phase of conversion or simplification from the tension peak, often marking the resolution of a struggle in which one side has attempted to generate complications. Consistent with established principles of high-level play, players with a decisive advantage tend to trade pieces to limit counterplay [37]. Our results confirm that this reduction is not the cause of the outcome, but rather its consequence: both humans and engines simplify positions once a win or a forced draw is already favored.
@@ II Result 2: Strategic tension reflects, does not predict game result
RESULT 3: Variability in strategic tension reflects different gameplay
While tension seemingly captures only general and coarse features of the game, it is correlated with key aspects of gameplay that distinguish humans from AI (see Fig. 9 in the Supplementary Materials).
@@
Number of links and loops: AI interaction networks exhibit more links and loops, indicating a richer web of mutual threats, defenses, and controls. However, the maximum average number of loops is higher in human games, reflecting the peak of human tension.
Number of pieces: AI maintains a higher number of pieces over the trajectory of games on average. This reflects a weaker tendency to simplify the game by removing pieces from the board. From mid to endgame, the small number of pieces limits the strategic tension that is typically reached in human games.
@@
Figure 3: Comparison of structural properties across human and AI games. We analyze the set of games shown in Fig. 2. Two measures are considered: the average standard deviation of the degrees associated with chess pieces, normalized by the mean of the degrees, and the average tension per piece. Red is for AI and blue for human games.
Degree distribution: The noise coefficient of the degree distribution over chess pieces (defined as the standard deviation over the mean degree) mirrors the rise and fall of strategic tension, first falling then rising. While human and AI games display a broadly similar global trend, human play exhibits more pronounced variation as shown in Fig. 3a. In particular, during the opening phase, the relative degree variation decreases markedly, reflecting a positioning phase in which interactions are more evenly distributed across pieces as tension accumulates. As the game progresses, however, the interaction network in human games exhibits greater variability, eventually returning to levels similar to those observed in AI games, or even lower.
@@
Tension per piece: As shown in Fig. 3b, the tension per piece $\lambda_{1}/n_{t}$ remains consistently higher across all plies in human games. While human play shows an approximately linear increase in tension per piece throughout the game, AI games display a sublinear progression that appears to plateau around ply 80. Human games involve more frequent captures of pieces (see Supplementary Materials), but the remaining pieces are then involved in a greater number and greater complexity of interactions.
Attack-defense balance: When distinguishing between attack and defense links, we find that the surplus of attack links over defense links, or “attack-defense balance” as shown in Fig. 4, moves inversely with the tension profile. It first drops in the early phase more defense-oriented phase of the game. As the game progresses, attack links gradually outnumber defense links. The curve is shifted towards more positive values in human games compared to AI games.
@@
Figure 4: Balance between attack and defense links across human and AI games. We analyze the set of games shown in Fig. 2. For both human and AI games, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively.
@@
Overall, we find that strategic tension aligns with basic gameplay metrics, such as the number of pieces and, in particular, the presence of loops in the interaction graph. At the same time, it is informative to relate it to the balance across pieces on the board and to how pieces are arranged in formations. Taken together, our results indicate that AI games retain more pieces, exhibit a more balanced network of attack–defense connections, and display less variability in the degree distribution across pieces.
@@
RESULT 4: Cumulative TENSION SCALES WITH SKILL AND COMPUTATIONAL COMPLEXITY
@@
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for rapid games (turquoise dots), blitz games (green dots), classical games played by top human players (blue star), and games between Stockfish and Leela from official TCEC tournaments (red star), all associated with estimated Elo ratings. Each dot/star represents the average across 1,200 games. The blue and red stars were derived from the data analyzed in Fig. 2. For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play. (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot. The two panels also show the $R^{2}$ values of the corresponding linear regressions.